OpenStudy (anonymous):

Find the variance of the sample data set. 127 135 128 131 133 127 122 A. 2.0 B. 4.4 C. 19.0 D. 4.0

OpenStudy (anonymous):

@rainbow.bubbles

OpenStudy (anonymous):

Order the values: 122, 127, 127, 128, 131, 133, 135

OpenStudy (anonymous):

i did that

OpenStudy (anonymous):

19 is your answer! :)

OpenStudy (anonymous):

can you explain how ?

OpenStudy (anonymous):

\[s^2= \frac{ 1 }{ n-1 } \sum_{i=1}^{n}(x _{i} - x)^2\]

OpenStudy (anonymous):

so just follow that equation ?